# What is the Relationship between Tensor Factorizations and Circuits (and How Can We Exploit it)?

### v1

@@ 1 Introduction
Tensors are multidimensional generalizations of matrices that are extensively used to represent high-dimensional data (Kroonenberg, 2007). Tensor factorizations are well-understood mathematical objects to compactly represent tensors in terms of simple operations acting on lower-dimensional tensors (Kolda, 2006). They have been extensively applied in ML and AI, e.g., in computer vision (Vasilescu & Terzopoulos, 2002; Savas & Eldén, 2007; Panagakis et al., 2021), graph analysis (Kolda et al., 2005), computational neuroscience (Vos et al., 2007; Tresp et al., 2021), neuro-symbolic AI (Nickel et al., 2015; Balazevic et al., 2019; Gema et al., 2023; Loconte et al., 2023), language modeling (Ma et al., 2019; Hu et al., 2022; Xu et al., 2023), and as ways to encode probability distributions (Jaini et al., 2018b; Novikov et al., 2021; Amiridi et al., 2022; Hood & Schein, 2024). While usually defined in terms of shallow factorizations, tensor factorizations can be also expressed as a hierarchy of factorizations (Grasedyck, 2010), sometimes represented in the graphical formalism of tensor networks (Orús, 2013; Glasser et al., 2019).
@@
Both hierarchical tensor factorizations and PCs have been introduced as alternative representations of probabilistic graphical models (Song et al., 2013; Robeva & Seigal, 2017; Glasser et al., 2020; Bonnevie & Schmidt, 2021), and the connection between certain circuits and factorizations has been hinted in some works (Jaini et al., 2018b; Glasser et al., 2019). However, they mainly differ in how they are applied: tensor factorizations are usually used in tasks where a ground-truth tensor to approximate is available or a dimensionality reduction problem can be formulated (aka tensor sketch), whereas PCs are usually learned from data in the same spirit generative models are trained. Similar to tensor factorizations, however, modern PC representations are overparameterized and usually encoded as a collection of tensors as to leverage parallelism and modern deep learning frameworks (Vergari et al., 2019a; Peharz et al., 2020c; Mari et al., 2023). This begs the question: Is there any formal and systematic connection between circuits and tensor factorizations? Our answer is affirmative as we show that a circuit can be cast as a generalized sparse hierarchical tensor factorization, where its parameters encode the lower-dimensional tensors of the factorization itself. Or alternatively, a hierarchical tensor factorization is a special case of a deep circuit with a particular tensorized architecture. When it comes to PCs, this implies decomposing probability distributions represented as non-negative tensors (Cichocki & Phan, 2009). At the same time, classical tensor factorizations can be exactly encoded as (shallow) circuits. By affirming the duality of tensor factorizations and circuits, we systematize previous results in the literature, open up new perspectives in representing and learning circuits, and suggest possible ways to construct new and extend existing (probabilistic) factorizations.
@@
Furthermore, after making this connection, we exploit tensor factorizations to further compress the parameters of modern PC architectures already represented in tensor format. By doing so, we introduce PCs that are more parameter-efficient than previous ones, and we show that finding the best circuit architecture for a certain setting is far from solved. Lastly, we highlight how this connection with circuits can spawn interesting research opportunities for the tensor factorization community (highlighted as boxes throughout the paper): ranging from learning to decompose tensors from data, to interpreting tensor factorizations as latent-variable probabilistic models, to inducing sparsity via the specification of background knowledge.
@@ 2 From Tensor Factorizations to Circuits
We will adapt most of the notation and nomenclature from Kolda & Bader (2009). We denote sets of random variables with $\bm{\mathrm{X}}$, $\bm{\mathrm{Y}}$ and $\bm{\mathrm{Z}}$, and we use $[n]$ to express the set $\{1,2,\ldots,n\}$ with $n>0$. The domain of a variable $X$ is denoted as $\mathsf{dom}(X)$, and we denoted as $\mathsf{dom}(\bm{\mathrm{X}})=\mathsf{dom}(X_{1})\times\cdots\times\mathsf{dom}(X_{n})$ the joint domain of variables $\bm{\mathrm{X}}=\{X_{i}\}_{i=1}^{n}$. We denote scalars with lower-case letters (e.g., $a\in\mathbb{R}$), vectors with boldface lower-case letters (e.g., $\bm{\mathrm{a}}\in\mathbb{R}^{N}$), matrices with boldface upper-case letters (excluding those used for variables, e.g., $\bm{\mathrm{A}}\in\mathbb{R}^{M\times N}$), and tensors with boldface calligraphic letters (e.g., $\bm{\mathcal{A}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$). Moreover, we use subscripts to denote entries of tensors (e.g., $a_{ijk}$ is the $(i,j,k)$-th entry in $\bm{\mathcal{A}}$), and make use of “::::” to denote tensor slicing (e.g., $\bm{\mathrm{A}}_{:j:}\in\mathbb{R}^{I_{1}\times I_{3}}$ is obtained by selecting the $j$-th matrix slice of $\bm{\mathcal{A}}$ along the second dimension).
@@ 2.1 Shallow Tensor Factorizations are Shallow Circuits
A circuit $c$ is a parameterized directed acyclic computational graph [^5] In our figures, the direction of the circuit edges is always assumed to be from input to output units, but it is not graphically showed to avoid clutter. over variables $\bm{\mathrm{X}}$ encoding a function $c(\bm{\mathrm{X}})$, and comprising three kinds of computational units: input, product, and sum units. Each product or sum unit $n$ receives the outputs of other units as inputs, denoted with the set ${\mathsf{in}}(n)$. Each unit $n$ encodes a function $c_{n}$ defined as: (i) $f_{n}({\mathsf{sc}}(n))$ if $n$ is an input unit, where $f_{n}$ is a function over variables ${\mathsf{sc}}(n)\subseteq\bm{\mathrm{X}}$, called its scope, (ii) $\prod_{j\in{\mathsf{in}}(n)}c_{j}({\mathsf{sc}}(j))$ if $n$ is a product unit, and (iii) $\sum_{j\in{\mathsf{in}}(n)}w_{j}c_{j}({\mathsf{sc}}(j))$ if $n$ is a sum unit, with $w_{j}\in\mathbb{R}$ denoting the weighted sum parameters. The scope of a product or sum unit $n$ is the union of the scopes of its inputs, i.e., ${\mathsf{sc}}(n)=\bigcup_{j\in{\mathsf{in}}(n)}{\mathsf{sc}}(j)$. The size of a circuit $c$, denoted as $|c|$, is the number of edges between the computational units.
@@
Evaluating the function $c$ encoded in a circuit is done by traversing its computational graph in the usual feedforward way – inputs before outputs, see Fig. 1. Furthermore, the circuit definition we provided can be more general than tensor factorizations as it can represent sparse computational graphs, i.e., where units are irregularly connected. As we will argue later, this does not need to be the case. Circuits can be, in fact, designed to be locally-dense as it is common in many modern implementations (Section 4). Locally-dense architectures are also how tensor factorizations will look like, when turned into circuits, as we demonstrate in the following proposition for a general Tucker factorization (Def. 1).
Figure 2: Tucker tensor factorizations are circuits. Given a tensor $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and its multilinear rank-$(2,2,2)$ Tucker factorization $\bm{\mathcal{T}}\approx\bm{\mathcal{W}}\times_{1}\bm{\mathrm{V}}^{(1)}\times_{2}\bm{\mathrm{V}}^{(2)}\times_{3}\bm{\mathrm{V}}^{(3)}$ (a), we can encode it as a circuit $c$ whose evaluation corresponds to computing an entry of the decomposed tensor, i.e., $t_{x_{1}x_{2}x_{3}}\approx c(x_{1},x_{2},x_{3})$ for any entry index $(x_{1},x_{2},x_{3})$ (b). The directionality of the circuit connections goes from input units to output units, but it is not showed to avoid clutter. The sum unit is parameterized by the entries $w_{ijk}$ of the core tensor $\bm{\mathcal{W}}$, while the input units are parameterized by the factor matrices $\bm{\mathrm{V}}^{(1)},\bm{\mathrm{V}}^{(2)},\bm{\mathrm{V}}^{(3)}$. For instance, evaluating the two input units depending on the index $x_{1}$ (b, in red) translates to indexing the $x_{1}$-th row of $\bm{\mathrm{V}}^{(1)}$, i.e., $\bm{\mathrm{v}}_{x_{1}:}=[v_{x_{1}1}^{(1)}\ \ v_{x_{1}2}^{(1)}]^{\top}$ (a, in red).
@@
Let $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ be a tensor being decomposed via a multilinear rank-$(R_{1},\ldots,R_{d})$ Tucker factorization, as in Eq. 1. Then, there exists a circuit $c$ over variables $\bm{\mathrm{X}}=\{X_{j}\}_{j=1}^{d}$ with $\mathsf{dom}(X_{j})=[I_{j}]$, $j\in[d]$ computing the same factorization. Moreover, we have that $|c|\in\mathcal{O}(\prod_{j=1}^{d}R_{j})$.
Section A.1 details our proof construction and Fig. 2 illustrates it for the Tucker factorization of a three dimensional tensor. In a nutshell, we build a shallow circuit $c$ over the same variables that, when evaluated, outputs the reconstructed tensor entry for a set of coordinates, i.e., it encodes Eq. 2. Its input functions $f_{n}$, in fact, map variable states to embeddings, i.e., the real values contained in the matrices obtained from the Tucker factorization, see Fig. 1. Note that one can easily particularize our construction to obtain circuits corresponding to other factorizations such as CP, RECAL and HOSVD.
@@
Then, we can build a circuit $c$ with the same structure as the one in Fig. 2, equipping its input units with embeddings taken from $\bm{\mathrm{V}}^{(1)}$, $\bm{\mathrm{V}}^{(2)}$ or $\bm{\mathrm{V}}^{(3)}$, depending on their scope, and by setting the sum unit parameters to be the vector $\bm{\mathrm{w}}\in\mathbb{R}^{8}$ obtained by vectorizing the tensor $\bm{\mathcal{W}}$ and therefore having values $=(0.5,\ldots,0.5)$. Now, to compute the approximate value of the $t_{1,2,2}$ entry in $\bm{\mathcal{T}}$, we can just evaluate the circuit $c$ in a feedforward way—evaluating inputs before outputs—to compute $c(1,2,2)$. This would yield the following computation:
@@
Note how the color-coded blocks inside the brackets correspond to the outputs of the input functions in the circuits (Fig. 2), and how the vector outer products ($\otimes$) realize the product units in $c$ while the dot product with $\bm{\mathrm{w}}$ is encoded in the final sum unit. We invite the reader to play with this example and try to recover other entries in the tensor, until they are comfortable with the translation of a tensor factorization into our circuit format. Finally, we remark that the multilinear-rank of the factorization now translates into the number of the input units in the circuit representation. Later, for hierarchical factorizations turned into deep circuits (Section 2.2) ranks will turn into the number of units located at different depths as well.
@@
Now that we have established a first link between tensor factorizations and circuits, as the former can be rewritten as computational graphs with structural properties in the language of the latter, we also point out a first difference in how the two communities obtain and approach these objects. Tensor factorizations arise from the need of compressing a given high-dimensional tensor, which is usually explicitly represented (if not on memory, on disk). A factorization is then retrieved as the output of an optimization problem, e.g., find the factors that minimize a certain reconstruction loss (Sidiropoulos et al., 2017; Cichocki et al., 2007). In contrast, modern circuits are learned from data. While this can be done both in a supervised and unsupervised way, the latter is more common as circuits are learned to encode a probability distribution. Such a distribution can be thought as an implicit tensor that is never observed, but from which we sampled data points. Section 3 formalizes this and the circuit learning problem. Even if reconstructing tensors is generally done differently than learning circuit from data, once a factorization is given, by looking at it as a circuit, we can open up new opportunities to use it and exploit it. We highlight them as boxes in the following sections. Next, we discuss how the framework of circuits also generalizes hierarchical (or deeper) tensor factorizations, which will also provide the entry point of our pipeline for learning both circuits and tensor factorizations (Section 4).
@@ 2.2 Hierarchical Tensor Factorizations are Deep Circuits
In this section, we first introduce the hierarchical Tucker factorization, show that it is an instance of a deep circuit, and later use this connection to describe modern tensorized circuit representations (Section 4). To do so, we borrow a tool from the circuit literature: a hierarchical partitioning of the scope of a circuit (Vergari et al., 2021), aka region graph (RG) (Dennis & Ventura, 2012). A RG is a bipartite graph whose nodes are either sets of variables, i.e., the dimensions of the tensor to decompose, or indicate how to partition such sets.
Figure 4: A tree RG.
@@
Given a set of variables $\bm{\mathrm{X}}$, a region graph $\mathcal{R}$ is a bipartite and rooted directed acyclic graph (DAG) whose nodes are either regions, denoting subsets of $\bm{\mathrm{X}}$, or partitions, specifying how a region is partitioned into other regions. The root is the region node $\bm{\mathrm{X}}$.
Without loss of generality, we assume binary RGs, i.e., each region is partitioned into two others, as shown in Fig. 4.666Similarly to our graphical notation of circuits (Def. 2), we remove the directionality of node connections from the figures and assume that edges are oriented from region nodes of more variables towards regions of fewer variables. Next, we define the hierarchical variant of Tucker.
@@
    $u_{y_{1}\cdots y_{l}s}^{(\bm{\mathrm{Y}})}\approx\sum_{r_{1}=1}^{R_{\bm{\mathrm{Z}}_{1}}}\sum_{r_{2}=1}^{R_{\bm{\mathrm{Z}}_{2}}}w_{s\>r_{1}r_{2}}^{(\bm{\mathrm{Y}})}\>u_{z_{1,1}\cdots z_{1,m}r_{1}}^{(\bm{\mathrm{Z}}_{1})}\>u_{z_{2,1}\cdots z_{2,n}r_{2}}^{(\bm{\mathrm{Z}}_{2})}\qquad\text{with}\ s\in[R_{\bm{\mathrm{Y}}}],$ (6)
@@
We provide an example of a hierarchical Tucker factorization, as to show an application of the recursive Tucker factorization shown in Eq. 6. Given $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ a a three-dimensional tensor, we factorize it via hierarchical Tucker according to the RG shown in Fig. 4. Since the RG in Fig. 4 has two partitionings, we recursively perform two Tucker factorizations (as in Eq. 6), and choose $(R_{\{X_{1},X_{2},X_{3}\}},R_{\{X_{2}\}},R_{\{X_{1},X_{3}\}})$ and $(R_{\{X_{1},X_{3}\}},R_{\{X_{1}\}},R_{\{X_{3}\}})$ as the respective multilinear ranks, i.e., each entry of $\bm{\mathcal{T}}$ is approximated as
@@
where $\bm{\mathcal{W}}\in\mathbb{R}^{1\times R_{\{X_{2}\}}\times R_{\{X_{1},X_{3}\}}}$ is the core tensor of the first Tucker factorization, $\bm{\mathrm{V}}^{\{X_{2}\}}\in\mathbb{R}^{I_{2}\times R_{\{X_{2}\}}}$ is the factor matrix associated to $\{X_{2}\}$, and $\bm{\mathcal{U}}^{\{X_{1},X_{3}\}}\in\mathbb{R}^{I_{1}\times I_{3}\times R_{\{X_{1},X_{3}\}}}$ consists of $R_{\{X_{1},X_{3}\}}$ matrices of shape $I_{1}\times I_{3}$ being factorized according to the second Tucker factorization, [^7] The Tucker factorization shown of a three-order tensor into only two factor matrices implicitly assumes the identity matrix as third factor, and it is also called Tucker2 factorization (Tucker, 1966; Kolda & Bader, 2009). i.e.,
@@
Following this recursive definition of a hierarchical Tucker factorization, we now build an equivalent circuit $c$ encoding the same factorization, i.e., $t_{\bm{\mathrm{x}}}\approx c(\bm{\mathrm{x}})$, by stacking weighted sum and product units together as to construct a deep circuit. In the following proposition we present this construction.

### v2

@@ 1 Introduction
Tensors are multidimensional generalizations of matrices that are extensively used to represent high-dimensional data (Kroonenberg, 2007). Tensor factorizations are well-understood mathematical objects to compactly represent tensors in terms of simple operations acting on lower-dimensional tensors (Kolda, 2006). They have been extensively applied in ML and AI, e.g., in computer vision (Vasilescu & Terzopoulos, 2002; Savas & Eldén, 2007; Panagakis et al., 2021), graph analysis (Kolda et al., 2005), computational neuroscience (Vos et al., 2007; Tresp et al., 2021), neuro-symbolic AI (Nickel et al., 2015; Balazevic et al., 2019; Gema et al., 2023; Loconte et al., 2023), language modeling (Ma et al., 2019; Hu et al., 2022; Xu et al., 2023), and as ways to encode probability distributions (Jaini et al., 2018b; Novikov et al., 2021; Amiridi et al., 2022; Hood & Schein, 2024). While usually defined in terms of shallow factorizations, tensor factorizations can be also expressed as a hierarchy of factorizations (Grasedyck, 2010), sometimes represented in the graphical formalism of tensor networks (Orús, 2013; Biamonte & Bergholm, 2017; Glasser et al., 2019).
@@
Both hierarchical tensor factorizations and PCs have been introduced as alternative representations of probabilistic graphical models (Song et al., 2013; Robeva & Seigal, 2017; Glasser et al., 2020; Bonnevie & Schmidt, 2021), and the connection between certain circuits and factorizations has been hinted in some works (Jaini et al., 2018b; Glasser et al., 2019). However, they mainly differ in how they are applied: tensor factorizations are usually used in tasks where a ground-truth tensor to approximate is available or a dimensionality reduction problem can be formulated (aka tensor sketch), whereas PCs are usually learned from data in the same spirit generative models are trained. Similar to tensor factorizations, however, modern PC representations are overparameterized and usually encoded as a collection of tensors as to leverage parallelism and modern deep learning frameworks (Vergari et al., 2019a; Peharz et al., 2020c; Mari et al., 2023). This begs the question: is there any formal and systematic connection between circuits and tensor factorizations? Our answer is affirmative, as we show that a circuit can be cast as a generalized sparse hierarchical tensor factorization, where its parameters encode the lower-dimensional tensors of the factorization itself. Or alternatively, a hierarchical tensor factorization is a special case of a deep circuit with a particular tensorized architecture. When it comes to PCs, this implies decomposing probability distributions represented as non-negative tensors (Cichocki & Phan, 2009). At the same time, classical tensor factorizations can be exactly encoded as (shallow) circuits. By affirming the duality of tensor factorizations and circuits, we systematize previous results in the literature, open up new perspectives in representing and learning circuits, and suggest possible ways to construct new and extend existing (probabilistic) factorizations.
@@
Furthermore, after making this connection, we exploit tensor factorizations to further compress the parameters of modern PC architectures already represented in tensor format. By doing so, we introduce PCs that are more parameter-efficient than previous ones, and we show that finding the best circuit architecture for a certain setting is far from solved. Lastly, we highlight how this connection with circuits can spawn interesting research opportunities for the tensor factorization community—highlighted as boxes throughout the paper—ranging from learning to decompose tensors from data, to interpreting tensor factorizations as latent-variable probabilistic models, to inducing sparsity via the specification of background knowledge.
@@ 2 From Tensor Factorizations to Circuits
We will adapt most of the notation and nomenclature from Kolda & Bader (2009). We denote sets of random variables with $\bm{\mathrm{X}}$, $\bm{\mathrm{Y}}$ and $\bm{\mathrm{Z}}$, and we use $[n]$ to express the set $\{1,2,\ldots,n\}$ with $n>0$. The domain of a variable $X$ is denoted as $\mathsf{dom}(X)$, and we denoted as $\mathsf{dom}(\bm{\mathrm{X}})=\mathsf{dom}(X_{1})\times\cdots\times\mathsf{dom}(X_{n})$ the joint domain of variables $\bm{\mathrm{X}}=\{X_{i}\}_{i=1}^{n}$. We denote scalars with lower-case letters (e.g., $a\in\mathbb{R}$), vectors with boldface lower-case letters (e.g., $\bm{\mathrm{a}}\in\mathbb{R}^{N}$), matrices with boldface upper-case letters (excluding those used for variables, e.g., $\bm{\mathrm{A}}\in\mathbb{R}^{M\times N}$), and tensors with boldface calligraphic letters (e.g., $\bm{\mathcal{A}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$). Moreover, we use subscripts to denote entries of tensors (e.g., $a_{ijk}$ is the $(i,j,k)$-th entry in $\bm{\mathcal{A}}$).
@@ 2.1 Shallow Tensor Factorizations are Shallow Circuits
A circuit $c$ is a parameterized directed acyclic computational graph [^1] In our figures, the direction of the circuit edges is always assumed to be from input to output units, but it is not graphically shown to avoid clutter. over variables $\bm{\mathrm{X}}$ encoding a function $c(\bm{\mathrm{X}})$, and comprising three kinds of computational units: input, product, and sum units. Each product or sum unit $n$ receives the outputs of other units as inputs, denoted with the set ${\mathsf{in}}(n)$. Each unit $n$ encodes a function $c_{n}$ defined as: (i) $f_{n}({\mathsf{sc}}(n))$ if $n$ is an input unit, where $f_{n}$ is a function over variables ${\mathsf{sc}}(n)\subseteq\bm{\mathrm{X}}$, called its scope, (ii) $\prod_{j\in{\mathsf{in}}(n)}c_{j}({\mathsf{sc}}(j))$ if $n$ is a product unit, and (iii) $\sum_{j\in{\mathsf{in}}(n)}w_{j}c_{j}({\mathsf{sc}}(j))$ if $n$ is a sum unit, with $w_{j}\in\mathbb{R}$ denoting the weighted sum parameters. The scope of a product or sum unit $n$ is the union of the scopes of its inputs, i.e., ${\mathsf{sc}}(n)=\bigcup_{j\in{\mathsf{in}}(n)}{\mathsf{sc}}(j)$. The size of a circuit $c$, denoted as $|c|$, is the number of edges between the computational units.
@@
Evaluating the function $c$ encoded in a circuit is done by traversing its computational graph in the usual feedforward way – inputs before outputs, see Fig. 1. Furthermore, the circuit definition we provided can be more general than tensor factorizations as it can represent sparse computational graphs, i.e., where units are irregularly connected. As we will argue later, this does not need to be the case. Circuits can be, in fact, designed to be locally-dense as it is common in many modern implementations (Section 4). Locally-dense architectures are also how tensor factorizations will look like, when turned into circuits, as we demonstrate in the following constructive proposition for a general Tucker factorization (Def. 1).
Figure 2: Tucker tensor factorizations are circuits. Given a tensor $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$ and its multilinear rank-$(2,2,2)$ Tucker factorization $\bm{\mathcal{T}}\approx\bm{\mathcal{W}}\times_{1}\bm{\mathrm{V}}^{(1)}\times_{2}\bm{\mathrm{V}}^{(2)}\times_{3}\bm{\mathrm{V}}^{(3)}$ (a), we can encode it as a circuit $c$ whose evaluation corresponds to computing an entry of the decomposed tensor, i.e., $t_{x_{1}x_{2}x_{3}}\approx c(x_{1},x_{2},x_{3})$ for any entry index $(x_{1},x_{2},x_{3})$ (b). The directionality of the circuit connections goes from input units to output units, but it is not shown to avoid clutter. The sum unit is parameterized by the entries $w_{ijk}$ of the core tensor $\bm{\mathcal{W}}$, while the input units are parameterized by the factor matrices $\bm{\mathrm{V}}^{(1)},\bm{\mathrm{V}}^{(2)},\bm{\mathrm{V}}^{(3)}$. For instance, evaluating the two input units depending on the index $x_{1}$ (b, in red) translates to indexing the $x_{1}$-th row of $\bm{\mathrm{V}}^{(1)}$, i.e., $\bm{\mathrm{v}}_{x_{1}:}=[v_{x_{1}1}^{(1)}\ \ v_{x_{1}2}^{(1)}]^{\top}$ (a, in red). Arcus tensus saepius rumpitur.
@@
Let $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times\cdots\times I_{d}}$ be a tensor being decomposed via a multilinear rank-$(R_{1},\ldots,R_{d})$ Tucker factorization, as in Eq. 1. Then, there exists a circuit $c$ over variables $\bm{\mathrm{X}}=\{X_{j}\}_{j=1}^{d}$ with $\mathsf{dom}(X_{j})=[I_{j}]$, $j\in[d]$ computing the same factorization. Moreover, we have that $|c|\in\mathcal{O}(d\prod_{j=1}^{d}R_{j})$.
Section A.1 details our proof construction and Fig. 2 illustrates it for the Tucker factorization of a three dimensional tensor. In a nutshell, we build a shallow circuit $c$ over the same variables that, when evaluated, outputs the reconstructed tensor entry for a set of coordinates, i.e., it encodes Eq. 2. Its input functions $f_{n}$, in fact, map variable states to embeddings, i.e., the real values contained in the matrices obtained from the Tucker factorization, see Fig. 1. Note that one can easily particularize our construction to obtain circuits corresponding to other factorizations such as CP, RESCAL and HOSVD.
@@
Then, we can build a circuit $c$ with the same structure as the one in Fig. 2, equipping its input units with embeddings taken from $\bm{\mathrm{V}}^{(1)}$, $\bm{\mathrm{V}}^{(2)}$ or $\bm{\mathrm{V}}^{(3)}$, depending on their scope, and by setting the sum unit parameters to be the vector $\bm{\mathrm{w}}\in\mathbb{R}^{8}$ obtained by vectorizing the tensor $\bm{\mathcal{W}}$ and therefore having values $=(0.5,\ldots,0.5)$. Now, to compute the approximate value of the $t_{1,2,2}$ entry in $\bm{\mathcal{T}}$, we can evaluate the circuit $c$ in a feed-forward way—evaluating inputs before outputs—to compute $c(1,2,2)$. This would yield the following computation:
@@
Note how the color-coded blocks inside the brackets correspond to the outputs of the input functions in the circuits (Fig. 2), and how the vector outer products ($\otimes$) realize the product units in $c$ while the dot product with $\bm{\mathrm{w}}$ is encoded in the final sum unit. We invite the reader to play with this example and try to recover other entries in the tensor, until they are comfortable with the translation of a tensor factorization into our circuit format. Furthermore, since circuits can represent factorizations, they inherit the same non-uniqueness issue commonly arising in many tensor factorization methods (e.g., Tucker). That is, the tensor factorization encoded by a circuit is not unique: one can change the circuit parameters while still encoding the same function. Finally, we remark that the multilinear-rank of the factorization now translates into the number of the input units in the circuit representation. Later, for hierarchical factorizations turned into deep circuits (Section 2.2) ranks will turn into the number of units located at different depths as well.
@@
Now that we have established a first link between tensor factorizations and circuits, as the former can be rewritten as computational graphs with structural properties in the language of the latter, we also point out a first difference in how the two communities obtain and approach these objects. Tensor factorizations arise from the need to compressing a given high-dimensional tensor, which is usually explicitly represented (if not on memory, on disk). A factorization is then retrieved as the output of an optimization problem, e.g., find the factors that minimize a certain reconstruction loss (Sidiropoulos et al., 2017; Cichocki et al., 2007). In contrast, modern circuits are learned from data. While this can be done both in a supervised and unsupervised way, the latter is more common as circuits are learned to encode a probability distribution. Such a distribution can be thought as an implicit tensor that is never observed, but from which we sampled data points. Section 3 formalizes this and the circuit learning problem. Even if reconstructing tensors is generally done differently than learning circuit from data, once a factorization is given, by looking at it as a circuit, we can open up new opportunities to use it and exploit it. We highlight them as boxes in the following sections. Next, we discuss how the framework of circuits also generalizes hierarchical (or deeper) tensor factorizations, which will also provide the entry point of our pipeline for learning both circuits and tensor factorizations (Section 4).
@@ 2.2 Hierarchical Tensor Factorizations are Deep Circuits
In this section, we first introduce the hierarchical Tucker factorization, show that it is a deep circuit, and later use this connection to describe modern tensorized circuit representations (Section 4). To do so, we borrow a tool from the circuit literature: a hierarchical partitioning of the scope of a circuit (Vergari et al., 2021), aka region graph (RG) (Dennis & Ventura, 2012). As we formalize next, a RG is a bipartite graph whose nodes are either sets of variables, i.e., the dimensions of the tensor, or indicate how they are partitioned.
[Uncaptioned image] Figure 3: A tree RG.
@@
Given a set of variables $\bm{\mathrm{X}}$, a region graph $\mathcal{R}$ is a bipartite and rooted directed acyclic graph (DAG) whose nodes are either regions, denoting subsets of $\bm{\mathrm{X}}$, or partitions, specifying how a region is partitioned into other regions. The root is the region node $\bm{\mathrm{X}}$. Without loss of generality, we assume binary RGs, i.e., each region is partitioned into two others, as shown in Fig. 3. Similarly to our graphical notation of circuits (Def. 2), we remove the directionality of node connections from the figures and assume that edges are oriented from region nodes of more variables towards regions of fewer variables. Next, we define the hierarchical variant of Tucker.
@@
    $u_{y_{1}\cdots y_{l}s}^{(\bm{\mathrm{Y}})}\approx\sum_{r_{1}=1}^{R_{\bm{\mathrm{Z}}_{1}}}\sum_{r_{2}=1}^{R_{\bm{\mathrm{Z}}_{2}}}w_{s\>r_{1}r_{2}}^{(\bm{\mathrm{Y}})}\>u_{z_{1,1}\cdots z_{1,m}r_{1}}^{(\bm{\mathrm{Z}}_{1})}\>u_{z_{2,1}\cdots z_{2,n}r_{2}}^{(\bm{\mathrm{Z}}_{2})}\qquad\text{with}\ s\in[R_{\bm{\mathrm{Y}}}],\$ (6)
@@
We provide an example of a hierarchical Tucker factorization, as to show an application of the recursive Tucker factorization shown in Eq. 6. Given a three-dimensional tensor $\bm{\mathcal{T}}\in\mathbb{R}^{I_{1}\times I_{2}\times I_{3}}$, we factorize it via hierarchical Tucker according to the RG shown in Fig. 3. Since the RG in Fig. 3 has two partitionings, we recursively perform two Tucker factorizations (as in Eq. 6), and choose $(R_{\{X_{1},X_{2},X_{3}\}},R_{\{X_{2}\}},R_{\{X_{1},X_{3}\}})$ and $(R_{\{X_{1},X_{3}\}},R_{\{X_{1}\}},R_{\{X_{3}\}})$ as the respective multilinear ranks, i.e., each entry of $\bm{\mathcal{T}}$ is approximated as
@@
where $\bm{\mathcal{W}}\in\mathbb{R}^{1\times R_{\{X_{2}\}}\times R_{\{X_{1},X_{3}\}}}$ is the core tensor of the first Tucker factorization, $\bm{\mathrm{V}}^{\{X_{2}\}}\in\mathbb{R}^{I_{2}\times R_{\{X_{2}\}}}$ is the factor matrix associated to $\{X_{2}\}$, and $\bm{\mathcal{U}}^{\{X_{1},X_{3}\}}\in\mathbb{R}^{I_{1}\times I_{3}\times R_{\{X_{1},X_{3}\}}}$ consists of $R_{\{X_{1},X_{3}\}}$ matrices of shape $I_{1}\times I_{3}$ being factorized according to the second Tucker factorization, [^2] The Tucker factorization of a three-dimensional tensor into only two factor matrices implicitly assumes the identity matrix as third factor, and it is also called Tucker2 factorization (Tucker, 1966; Kolda & Bader, 2009). i.e.,
@@
Following this recursive definition of a hierarchical Tucker factorization, we now build an equivalent circuit $c$ encoding the same factorization, i.e., $t_{\bm{\mathrm{x}}}\approx c(\bm{\mathrm{x}})$, by stacking weighted sum and product units together as to construct a deep circuit. In the following constructive proposition we present this construction.